# ST-Booster: An Iterative SpatioTemporal Perception Booster for Vision-and-Language Navigation in Continuous Environments

### v1

@@ IV-E Influence of Design Parameters
[…]
GAH Weights. As a coarse-grained expert prior, GAH is applied by multiplying it with a weight and adding it to the probability heatmap predicted by the original waypoint predictor. This experiment analyzes the sensitivity of performance on val unseen split dataset to different GAH weights. We tested three weight values, as shown in Table IV. Reducing the GAH weight from $10^{-4}$ to $10^{-5}$ improves most metrics. However, further reduction to $10^{-6}$ slightly degrades performance. This indicates that while GAH provides useful prior knowledge, its coarse-grained nature can introduce bias limit exploration, and reduce performance. In this paper, we use a GAH weight of $10^{-5}$ with the best performance for the remaining experiments.
TABLE III: The effect of grid map size on navigation performance
[…]
TABLE IV: The effect of GAH weight on navigation performance
[…]
$1.0\times 10^{-5}$ 13.14 4.77 67.6 61.0 50.2
[…]
@@ V Conclusion
To tackle the perception challenges in VLN-CE, this paper proposes an iterative spatiotemporal enhancement method — the ST-Booster, which encompasses three interconnected modules. The hierarchical spatiotemporal encoding (HSTE) module decouples environmental understanding into long-term navigation memory encoded in topological graphs and short-term geometric details captured by grid maps. The multi-granularity alignment fusion (MGAF) module introduces the bidirectional geometric-aware transformation to enable semantic-aware knowledge exchange across spatiotemporal scales. The value-guided waypoint generation (VGWG) module bridges perceptual features with actionable navigation priors. Extensive comparative experiments and comprehensive performance analysis are conducted, the results of which demonstrate that our proposed method outperforms the state-of-the-art approaches, particularly in complex tasks with environmental disturbances. Future work will extend this paradigm to tackle open-world perception challenges, including partial observability and lifelong map maintenance, further bridging the gap between virtual benchmarks and real-world deployment.

### v2

@@ IV-E Influence of Design Parameters
Grid Map Size. We analyze the effect of grid map size on navigation performance in the val unseen split of VLN-CE. We evaluate grid maps with sizes $\times7$, $\times11$, $\times15$} and cell resolutions of $0.5\,\text{m}$, $1\,\text{m}$} for grid maps. As illustrated in Table IV, increasing the cell size generally improves navigation performance. This improvement is caused by the fact that smaller cells capture more noise and redundant details, interfering with the agent’s extraction of key environmental cues. Furthermore, when the cell resolution is $0.5\,\text{m}$, increasing the map size boosts success rates by incorporating more environmental information. However, with a $1\,\text{m}$ cell resolution, the performance of the $\times15$ map declines. This is due to the excessive map size, which results in an increased region overlap between adjacent time steps. This causes redundant computations and data processing complexity. The results indicate that balancing scale size and cell resolution is crucial for optimizing navigation.
TABLE V: Performance comparison on the val-unseen dataset between random noised maps (RNMs) and GAHs under different map weight settings
Map GAH Weight TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$)
RNM $1.0\times 10^{-4}$ 13.12 5.0 65.3 56.9 47.2
$1.0\times 10^{-5}$ 13.13 4.8 66.7 58.4 48.6
$1.0\times 10^{-6}$ 13.18 4.9 66.9 59.6 49.6
GAH $1.0\times 10^{-4}$ 13.48 5.0 65.2 57.7 47.2
$1.0\times 10^{-5}$ 13.14 4.8 67.6 61.0 50.2
$1.0\times 10^{-6}$ 13.13 4.8 67.1 59.8 49.6
GAH Weights. We evaluate the effect of different GAH fusion weights to assess the contribution of GAH to navigation reasoning. As a coarse-grained expert prior, the GAH is fused with the predicted waypoint heatmap $\mathbf{P}_{t}$ via weighted addition $(\delta)$. The weight selection is first driven by the magnitude gap between $\mathbf{H}_{t}$ and $\mathbf{P}_{t}$. The GAH follows the scale of the ground-truth map $\mathbf{H}^{*}_{t}$, obtained by converting the next waypoint into a Gaussian-blurred heatmap in $[0,1]$ and applying a scaling factor $\rho=10$: $\mathbf{H}^{*}_{t}=\rho\cdot\frac{g_{i}}{g^{\max}}$. As a result, $\max{\mathbf{H}_{t}}=10$. In contrast, $\mathbf{P}_{t}$ is a softmax-normalized distribution with a total sum of $1$ and an average of approximately $7\times 10^{-4}$. This substantial scale mismatch requires using a small fusion weight for $\mathbf{H}_{t}$. Moreover, large weights make the fused waypoint distribution overly concentrated, amplifying prediction errors from $\mathbf{H}_{t}$ and degrading performance. Consistent with this analysis, experiments show that reducing the GAH weight from $10^{-4}$ to $10^{-5}$ yields performance improvements across most metrics, whereas further decreasing it to $10^{-6}$ results in a slight drop.
We compare GAH $\mathbf{H}_{t}$ with a Random Noise Map (RNM) having the same order of magnitude. At both $10^{-4}$ and $10^{-5}$, RNM performs worse than $\mathbf{H}_{t}$, with a 2.6% lower success rate at $10^{-5}$. This indicates that RNM changes waypoint distributions but fails to guide sampling towards task-relevant regions, while $\mathbf{H}_{t}$ effectively highlights relevant areas, improving waypoint inference. When the weight is reduced to $10^{-6}$, the performance gap narrows, showing that such a small weight has minimal impact on waypoint distribution. Overall, an appropriate weight provides balanced prior knowledge, while an excessively large weight may introduce bias and limit exploration. Based on these results, we adopt a GAH weight of $10^{-5}$ for optimal performance in subsequent experiments.
@@ V Conclusion
To tackle the perception challenges in VLN-CE, this paper proposes an iterative spatiotemporal enhancement method named ST-Booster, which encompasses three interconnected modules. The hierarchical spatiotemporal encoding (HSTE) module decouples environmental understanding into long-term navigation memory encoded in topological graphs and short-term geometric details captured by grid maps. The multi-granularity alignment fusion (MGAF) module introduces the bidirectional geometric-aware transformation to enable semantic-aware knowledge exchange across spatiotemporal scales. The value-guided waypoint generation (VGWG) module bridges perceptual features with actionable navigation priors. Extensive comparative experiments and comprehensive performance analysis are conducted, demonstrating that our proposed method outperforms the state-of-the-art approaches, particularly in complex tasks with environmental disturbances. Future work will extend this paradigm to tackle open-world perception challenges, including partial observability and lifelong map maintenance, further bridging the gap between virtual benchmarks and real-world deployment.